Any two groups of order m are isomorphic, where m ∈ N . True or False. justify
1
Expert's answer
2020-06-22T18:15:14-0400

Let’s consider two groups "Z_4" and "U(8)"

"Z_4=\{0,1,2,3\}" addition "\mod 4"

"U(8)=\{1,3,5,7\}" multiplication "\mod 8"

"|Z_4|=|U(8)|=4"

Suppose, that there exists isomorphism "f: Z_4\rightarrow U(8)" .

Then "f" will map the identity element of "Z_4"  to the identity element of "U(8)" : "f(0)=1" .

"f(1)" can be equal to "3,5,7."

"f(2)=f(1+1)=f(1)\times f(1)=1," because "3^2=5^2=7^2=1\mod 8" .

We have, that "f(2)=1," but "2" isn’t identity element of "Z_4" .

Contradiction.

So, there is no isomorphism between "Z_4" and "U(8)" .


Although both these groups have order 4, they aren’t isomorphic.


Answer: False.